determine whether or not the set of vectors {(1,2,-1),(0,3,1),(1,-5,3)} is a basis of R^3 also find dimension


1
Expert's answer
2021-06-14T15:36:29-0400

We can set up a matrix and use Gaussian elimination .


"\begin{pmatrix}
 1 & 2 & -1 \\
 0 & 3 & 1 \\
 1 & -5 & 3 \\
\end{pmatrix}"

"R_3=R_3-R_1"

"\begin{pmatrix}
 1 & 2 & -1 \\
 0 & 3 & 1 \\
 0 & -7 & 4 \\
\end{pmatrix}"

"R_2=R_2\/3"

"\begin{pmatrix}
 1 & 2 & -1 \\
 0 & 1 & 1/3 \\
 0 & -7 & 4 \\
\end{pmatrix}"

"R_1=R_1-2R_2"

"\begin{pmatrix}
 1 & 0 & -5/3 \\
 0 & 1 & 1/3 \\
 0 & -7 & 4 \\
\end{pmatrix}"

"R_3=R_3+7R_2"


"\begin{pmatrix}
 1 & 0 & -5/3 \\
 0 & 1 & 1/3 \\
 0 & 0 & 19/3\\
\end{pmatrix}"

"R_3=(3\/19)R_3"


"\begin{pmatrix}
 1 & 0 & -5/3 \\
 0 & 1 & 1/3 \\
 0 & 0 & 1\\
\end{pmatrix}"

"R_1=R_1+(5\/3)R_3"


"\begin{pmatrix}
 1 & 0 & 0 \\
 0 & 1 & 1/3 \\
 0 & 0 & 1\\
\end{pmatrix}"

"R_2=R_2-R_3\/3"


"\begin{pmatrix}
 1 & 0 & 0 \\
 0 & 1 & 0 \\
 0 & 0 & 1\\
\end{pmatrix}"

Since the rank of the matrix is 3, then the set {(1,2,-1),(0,3,1),(1,-5,3)} is a basis of "R^3."

The dimension is 3.